OpenStudy (asapbleh):

Find the volume between the surfaces given by: z=f(x,y)=2+2x-4y and z=g(x,y)=3+x^2 over the triangular plane region D with vertices (0,0), (3,0), and (3,6). { double or triple integral}

ganeshie8 (ganeshie8):

what have you tried so far

OpenStudy (asapbleh):

I tried graphing the plane and parabola. I have trouble graphing the plane and setting up the integral. Mainly the boundaries

OpenStudy (asapbleh):

@ganeshie8

OpenStudy (asapbleh):

I tried finding the axes for the plane. I got z=2, y=1/2, and for x, i sometimes get 1 and other times i get -1

OpenStudy (asapbleh):

but the top surface is 2+2z-4y and the bottom is 3+x^2 right?

ganeshie8 (ganeshie8):

it looks like this i think : http://gyazo.com/1f2ba93b9f44e8bae56d6e4d509782cf

OpenStudy (asapbleh):

ok so thats the graph i got the first time i did it, but i now im confused as to whos the top and whos the bottom

OpenStudy (asapbleh):

yay ur finally back

OpenStudy (asapbleh):

i just need clarification on whos the top surface and bottom

ganeshie8 (ganeshie8):

hey are u really sure there are no typoes in the question because we need to setup 3 integrals to find the volume :O

OpenStudy (asapbleh):

i triple checked. it could be either double or triple cus when we set it up as triple then the function inside is just 1 or a constant.

ganeshie8 (ganeshie8):

we need to setup 3 triple integrals

OpenStudy (asapbleh):

WHAT

OpenStudy (asapbleh):

not just 1 triple integral, but 3?

ganeshie8 (ganeshie8):

depends, if u want signed volume then 1 triple integral will do

ganeshie8 (ganeshie8):

lets assume they want the signed volume

OpenStudy (asapbleh):

ok i think its just 1 triple integral. you got me freaking out

ganeshie8 (ganeshie8):

\[\int\limits_0^3\int\limits_0^{2x} ((2+2x-4y) - (3+x^2)) ~dydx\] ?

OpenStudy (asapbleh):

Ok, so the plane is the top surface and the parabola is the bottom?

OpenStudy (asapbleh):

how can you tell? thats why im asking

ganeshie8 (ganeshie8):

its not same over the entire region, look at the earlier plot

OpenStudy (asapbleh):

i saw, but its like when the plane slices the parabola thingy, the values of x are negative, so i was hesistant to do that

ganeshie8 (ganeshie8):

i get what you mean, you need to setup 3 triple integrals if you want the actual volume

OpenStudy (asapbleh):

and what do you mean by not the same, i think our teacher just wants the main one, like the one nearest to the "origin"

ganeshie8 (ganeshie8):

if you want signed volume, just take the difference of functions

OpenStudy (asapbleh):

ahhhh so for the boundaries can u briefly tell me how u got them

ganeshie8 (ganeshie8):

look at the given region in xy plane

OpenStudy (asapbleh):

oh i see the 2x now. its the slope ahhhhhh omg i get this

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

yes :) do u have the final answer ?

OpenStudy (asapbleh):

so i get how to set up the integral now, so can you explain to me how to graph the plane. i keep on getting x=1 and x=-1

OpenStudy (asapbleh):

like when i put x=y=0, x=z=0, and y=z=0

ganeshie8 (ganeshie8):

z = 2+2x-4y find the intercepts

ganeshie8 (ganeshie8):

x intercept = -1 y intercept = 1/2 z intercept = 2 yes ?

OpenStudy (asapbleh):

yes i got that, but when you plug it back those intercepts back into the eqn, they're not equal

OpenStudy (asapbleh):

2=2+2(-1)-4(1/2) 2=2-2-2 2=/=-2

ganeshie8 (ganeshie8):

x intercept = -1, so the point would be `(-1, 0, 0)` y intercept = 1/2 z intercept = 2

ganeshie8 (ganeshie8):

you need to plugin the "point" corresponding to x intercept

ganeshie8 (ganeshie8):

do you know how to plot the line `x+y = 1` in xy plane ?

OpenStudy (asapbleh):

oh...right yeah. thanks for the reminder.

ganeshie8 (ganeshie8):

x intercept = 1 y intercept = 1 do u expect that line to satisfy the point (1, 1) ?

OpenStudy (asapbleh):

thanks for the help. appreciated

ganeshie8 (ganeshie8):

np :)